# Implementing Cumulative Functions with Generalized Cumulative Constraints

### v1

@@ Abstract
Modeling scheduling problems with conditional time intervals and cumulative functions has become a common approach when using modern commercial constraint programming solvers. This paradigm enables the modeling of a wide range of scheduling problems, including those involving producers and consumers. However, it is unavailable in existing open-source solvers and practical implementation details remain undocumented. In this work, we present an implementation of this modeling approach using a single, generic global constraint called the Generalized Cumulative. We also introduce a novel time-table filtering algorithm designed to handle tasks defined on conditional time-intervals. Experimental results demonstrate that this approach, combined with the new filtering algorithm, performs competitively with existing solvers enabling the modeling of producer and consumer scheduling problems and effectively scales to large problems.
Introduction
@@
The success of Constraint Programming as a modeling and solving technology for hard scheduling problems is well established (Laborie et al. 2018). The modeling has been further enhanced with the introduction of conditional time-interval variables (Laborie and Rogerie 2008), which represent the execution of tasks that may or may not be executed. Such variables allow to conveniently model problems with alternative resources or optional tasks such as in (Kinable, Wauters, and Berghe 2014; Kumar et al. 2018; Cappart and Schaus 2017; Kizilay, Eliiyi, and Van Hentenryck 2018). Building on conditional time-intervals, the same authors introduced an algebraical model for cumulative resources called cumulative functions (Laborie et al. 2009). It provides a convenient way to model producer-consumer problems with optional tasks such as in (Liu, Aleman, and Beck 2018; Gedik et al. 2018; Cappart et al. 2018; Thomas and Schaus 2024). CPOptimizer (Laborie et al. 2018) and OptalCP (Vilím and Pons 2023) are the only two solvers that fully support conditional time-interval variables and cumulative functions and this modeling paradigm is not supported by existing open-source solvers [^1] API languages such as Minizinc also support optional variables at the modeling layer (Mears et al. 2014).
@@
We propose an implementation of this modeling paradigm with lower and upper-bound restrictions on cumulative functions using a generalized cumulative constraint that supports negative heights (Beldiceanu and Carlsson 2002). This more general form of the cumulative constraint allows variable and negative resource consumptions as well as optional tasks. We introduce a new filtering algorithm for this constraint that is based on time-tabling (Gay, Hartert, and Schaus 2015a; Fahimi, Ouellet, and Quimper 2018; Letort, Carlsson, and Beldiceanu 2015) and does additional filtering on the height and length of tasks. The algorithm proceeds in two steps. First, it uses the Profile data structure recently introduced by (Gingras and Quimper 2016) and builds the optimistic/pessimistic profiles of resources produced/consumed by tasks. It then processes each task to prune its bounds by comparing it to the profile. The experimental results show that the newly introduced algorithm performs well compared to existing ones, particularly in large-scale scheduling problems solved with a greedy search.
Modeling with Cumulative Functions
Conditional Time-Intervals
@@
The domain of a conditional time-interval variable can be expressed as a subset of {}$\{\bot\}\cup\{[s,e)\mid s\leq e,s,e\in\mathbb{Z}\}$ (Laborie and Rogerie 2008). It is fixed if $x=\bot$ (the interval is not executed) or $x=[s,e)$ (the interval starts at $s$ and ends at $e$). For convenience, the description of the filtering algorithm manipulates conditional time-interval domains through the following attributes for each task $i\in T$:
@@
  • $s_{i}\in[\underline{s}_{i},\overline{s}_{i}]$ and $e_{i}\in[\underline{e}_{i},\overline{e}_{i}]$ are the start/end variables.
@@
A bound consistent filtering on the relation $s_{i}+d_{i}=e_{i}$ is maintained atomically on the corresponding attributes, and the domain of a time-interval variable becomes $\bot$ whenever one attribute range ($s,d$ or $e$) becomes empty which may trigger an inconsistency if $p_{i}=true$. The whole domain of a conditional time-interval $i\in T$ is thus encoded by a tuple of variables $x_{i}=\langle s_{i},d_{i},e_{i},p_{i}\rangle$.
Cumulative Functions
For modeling cumulative problems, (Laborie et al. 2009) introduced the idea of cumulative function expressions. They allow the expression of step-wise integer functions of conditional time-interval variables. The contribution of a single time-interval variable $i\in T$ is represented by an elementary cumulative function receiving an interval $x_{i}$ and a non-negative consumption value $c_{i}$ or a non-negative height range $[\underline{c}_{i},\overline{c}_{i}]$. The two main [^2] Variants exist such as stepAtEnd. elementary functions are pulse and stepAtStart (Laborie et al. 2009) as represented in Figure 1.
Figure 1: Elementary cumulative functions.
@@
Three tasks $A$, $B$ and $C$ contribute to a cumulative function $f=stepAtStart(A,2)\allowbreak-\allowbreak(pulse(B,1)\allowbreak+\allowbreak stepAtEnd(C,1))$. The resulting AST and cumulative function are represented in Figure 2. On the right, the solid parts represent the time windows of the tasks while the hatched parts represent their contribution to the profile. A dotted line represents the final cumulative function $f$.
Figure 2: Example 1: AST (left) and cumul. function (right).
A cumulative function can be flattened using Algorithm 1 to extract a set of cumulative tasks (tasks with their positive or negative heights). This algorithm traverses the AST and collects the activities at the leaf nodes, reversing the height signs for the activities flattened on the right branch of a minus node. Note that whenever encountering a $stepAtStart(x,c)$ leaf node, the cumulative task collected is defined on a fresh task interval $x^{\prime}$ related to the original one with: $x^{\prime}.s=x.s$ (same start as $x$), $x^{\prime}.p=x.p$ (same execution status), and $x^{\prime}.e=Horizon$ (ending at the time horizon) to reflect the definition of a step function [^3] Other cases such as $stepAtEnd(x,c)$ are defined similarly..
1 if $n=\text{Pulse}(x,c)$ then  $\{(x,c)\}$
2 else if $n=\text{StepAtStart}(x,c)$ then  $\{(x^{\prime},c)\}$
3 else if $n=\text{Plus}(n_{l},n_{r})$ then
4 $\text{flatten}(n_{l})\cup\text{flatten}(n_{r})$
5else if $n=\text{Minus}(n_{l},n_{r})$ then
6 $\text{flatten}(n_{l})\cup\{(x,-c)\mid(x,c)\in\text{flatten}(n_{r})\}$
7else  $error$ // Unknown pattern
[…]
@@ 6 $\text{flatten}(n_{l})\cup\{(x,-c)\mid(x,c)\in\text{flatten}(n_{r})\}$
Continuing example 1, $f$) collects the cumulative tasks $\{(A^{\prime},2),(B,-1),(C^{\prime},-1)\}$ illustrated on the right of Figure 2.
Constraints on Cumulative Functions
The height of a cumulative function can be restricted with a lower and upper bound at every time overlapping at least one executing time-interval: $\underline{C}\leq f\leq\overline{C}$. This constraint is called $alwaysIn(f,\underline{C},\overline{C})$ in (Laborie et al. 2009). It can be enforced by posting a GeneralizedCumulative constraint requiring a resource consumption between $\underline{C}$ and $\overline{C}$ for the set of cumulative tasks $T$ collected with the Flatten function. The whole domain of a conditional (cumulative) task $i\in T$ is encoded by a tuple of variables $\langle s_{i},d_{i},e_{i},c_{i},p_{i}\rangle$ where the integer variable $c_{i}$ represents its consumption (possibly negative or hybrid) on the resource. The set of tasks $T$ is partitioned into three subsets as given in Definition 1.
@@
The set $O=\{i\in T\mid p_{i}=\{true,false\}\}$ denotes the set of optional tasks, those for which it is not decided yet if they will execute or not on the resource. The set $R=\{i\in T\mid p_{i}=true\}$ denotes the set of required tasks, those which will surely be executed on the resource. The set $E=\{i\in T\mid p_{i}=false\}$ denotes the set of excluded tasks, those which will surely not be executed on the resource. The sets $O$, $R$ and $E$ are disjoint and form a partition of the set $T$, i.e., $T=O\cup R\cup E$.
@@
[…]
$\sum\limits_{i\in R\mid s_{i}\leq\tau<e_{i}}c_{i}\in[\underline{C},\overline{C}]\qquad\forall\tau\in\bigcup_{i\in R}[s_{i},e_{i})$ (1)
@@
Enforcing global bound-consistency for the cumulative constraint is NP-Hard (Garey and Johnson 1979). However, enforcing it on the decomposition into sum constraints at every point in the horizon, as in (1), is not NP-Hard. Unfortunately, this approach is inefficient for large horizons, as it requires a sum constraint to be enforced at every time point. Therefore, we aim to achieve the same filtering in polynomial time relative to the number of tasks, rather than the size of the horizon. This filtering is generally referred to as time-tabling filtering. Intuitively, it works by aggregating all the fixed parts (when $\underline{s}_{i}<\overline{e}_{i}$) to postpone task execution.
Timetabling Algorithm
The Profile Data Structure
The Profile data structure is an aggregation of adjoined rectangles of different lengths and heights introduced in (Gingras and Quimper 2016; Gay, Hartert, and Schaus 2015a) to record the resource utilization of tasks. The end of a rectangle is the beginning of the next one. The starting (resp. ending) of a rectangle is represented by a tuple called time point with a component representing the beginning (resp. ending) time of the rectangle and other components that contain information relative to the resource consumption for the duration of the rectangle. These time points are sorted in increasing order of time and are kept in a linked list structure called Profile. An original idea of (Gingras and Quimper 2016) is to have pointers $tp(\underline{s}_{i})$, $tp(\underline{e}_{i})$ and $tp(\overline{e}_{i})$ linking to the time point associated with $\underline{s}_{i}$, $\underline{e}_{i}$, and $\overline{e}_{i}$ respectively.
Timetable Filtering
[…]
@@
$\underline{P}_{t}=\sum\limits_{\mathclap{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}}\min(\underline{c}_{i},0)+\sum\limits_{\mathclap{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}}\max(\underline{c}_{i},0)$ (2)
$\overline{P}_{t}=\sum\limits_{\mathclap{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}}\max(\overline{c}_{i},0)+\sum\limits_{\mathclap{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}}\min(\overline{c}_{i},0)$ (3)
@@
Timeline Initialization and Consistency Check
The function initializeTimeline (Algorithm 2) initializes the timeline, links the corresponding time points to tasks, and computes the minimum and maximum profile of the resource utilization. It receives as input the set of tasks $T$. First, events are generated by iterating over the tasks that are not absent (Line 2). For each non-absent task $i$, two events are generated which correspond to the minimum start time ($\underline{s}_{i}$), and the maximum end time ($\overline{e}_{i}$) of the task. For tasks that are present and with a fixed part, two additional events are generated for the maximum start time ($\overline{s}_{i}$) and the minimum end time ($\underline{e}_{i}$) of the task.
1 ${}events\leftarrow\{\}$
@@
3 $events\leftarrow events\cup\{\underline{s}_{i},\overline{e}_{i}\}$
@@
5    $events\leftarrow events\cup\{\overline{s}_{i},\underline{e}_{i}\}$
[…]
7 $tp\leftarrow new\,tp(\min(events.time))$
8 $P\leftarrow\{tp\}$
9 for $e\in events$ sorted by time do
@@ 8 $P\leftarrow\{tp\}$
11 if $tp.time<e.time$ then
12    if #$\#fp>0\wedge(\underline{P}>\overline{C}\vee\overline{P}<\underline{C})$ then
13       return Fail
@@
15    $tp\leftarrow new\,tp(e.time)$
16    $P.append(tp)$
17
18 if $e=\underline{s}_{i}$ then
19    $\underline{P}\leftarrow\underline{P}+\min(\underline{c}_{i},0)$
20    $\overline{P}\leftarrow\overline{P}+\max(\overline{c}_{i},0)$
21    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
22    $tp(\underline{s}_{i})\leftarrow tp$
@@
24 if $e=\overline{e}_{i}$ then
25    $\underline{P}\leftarrow\underline{P}-\min(\underline{c}_{i},0)$
26    $\overline{P}\leftarrow\overline{P}-\max(\overline{c}_{i},0)$
27    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
28    $tp(\overline{e}_{i})\leftarrow tp$
29
30 if $e=\overline{s}_{i}$ then
31    $\underline{P}\leftarrow\underline{P}+\max(\underline{c}_{i},0)$
32    $\overline{P}\leftarrow\overline{P}+\min(\overline{c}_{i},0)$
33    #$\#fp\leftarrow\#fp+1$
34    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
35
36 if $e=\underline{e}_{i}$ then
37    $\underline{P}\leftarrow\underline{P}-\max(\underline{c}_{i},0)$
38    $\overline{P}\leftarrow\overline{P}-\min(\overline{c}_{i},0)$
39    #$\#fp\leftarrow\#fp-1$
40    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
41
@@ 38    $\overline{P}\leftarrow\overline{P}-\min(\overline{c}_{i},0)$
[…]
As the timeline is initialized, the consistency of the minimum and maximum profiles is checked for each time point at Line 2. Once a time point is completed (when creating the next time point), we ensure that if at least one task is guaranteed to execute during the time point (#$\#fp>0$), the minimum profile $\underline{P}$ is not above the maximum capacity of the resource $\overline{C}$ and the maximum profile $\overline{P}$ is not below the minimum capacity of the resource $\underline{C}$.
@@
$A$
: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=[1,2],\allowbreak p=true\rangle$;
$B$
: $\langle s=[2,4],\allowbreak d=[3,4],\allowbreak e=[5,7],\allowbreak c=2,\allowbreak p=true\rangle$;
$C$
: $\langle s=[3,8],\allowbreak d=[1,3],\allowbreak e=[4,9],\allowbreak c=[-2,1],\allowbreak p=\{true,false\}\rangle$;
and a resource $C$ of capacity bounds [0,1][0,1][ 0 , 1 ]. The time points and their attributes are shown in Table 1. Figure 3 shows a representation of the tasks and the resulting profile ranges. Note that the events $\overline{s}_{c}$ and $\underline{e}_{c}$ are not considered as task $C$ has no fixed part.
[…]

### v2

@@ Abstract.
Modeling scheduling problems with conditional time intervals and cumulative functions has become a common approach when using modern commercial constraint programming solvers. This paradigm enables the modeling of a wide range of scheduling problems, including those involving producers and consumers. However, it is unavailable in existing open-source solvers and practical implementation details remain undocumented. In this work, we present an implementation of this modeling approach using a single, generic global constraint called the Generalized Cumulative. We also introduce a novel time-table filtering algorithm specifically designed to handle tasks defined on conditional time-intervals. Experimental results demonstrate that this approach, combined with the new filtering algorithm, performs competitively with existing solvers enabling the modeling of producer and consumer scheduling problems and effectively scales to large-scale problems.
1. Introduction
@@
The success of Constraint Programming as a modeling and solving technology for hard scheduling problems is well established (laborie2018ibm). The modeling flexibility has been further enhanced with the introduction of conditional time-interval variables (LaborieR08), which represent the execution of tasks that may or may not be executed. Such variables allow to conveniently model problems with alternative resources or optional tasks such as in (kinable2014concrete; kumar2018station; cappart2017rescheduling; kizilay2018constraint). Building on top of conditional time-intervals, the same authors introduced an algebraical model for cumulative resources called cumulative functions (LaborieRSV09). It provides an efficient and convenient way to model producer-consumer problems with optional tasks such as in (liu2018modelling; gedik2018constraint; cappart2018constraint; thomas2024constraint). To the best of our knowledge, CPOptimizer (laborie2018ibm) and OptalCP (OptalCP) are the only two (commercial) solvers that fully support conditional time-interval variables and cumulative functions and this modeling paradigm is not supported by existing open-source solvers [^1] API languages such as Minizinc also support optional variables at the modeling layer (optionminizinc).
@@
We propose an implementation of this modeling paradigm with lower and upper-bound restrictions on cumulative functions using a generalized cumulative constraint that supports negative heights (BeldiceanuC02). This more general form of the cumulative constraint allows variable and negative resource consumptions as well as optional tasks. We introduce a new filtering algorithm for this constraint that is based on time-tabling (GayHS15; FahimiOQ18; LetortCB15) and does additional filtering on the height and length of tasks. The algorithm proceeds in two steps. First, it uses the Profile data structure recently introduced by (GingrasQ16) and builds the optimistic/pessimistic profiles of resources produced/consumed by tasks. It then processes each task to prune its bounds by comparing it to the profile. The experimental results show that the newly introduced algorithm performs well compared to existing ones, particularly in large-scale scheduling problems solved with a greedy search.
2. Modeling with Cumulative Functions
2.1. Conditional Time-Intervals
@@
The domain of a conditional time-interval variable can be expressed as a subset of $\{\bot\}\cup\{[s,e)\mid s\leq e,s,e\in\mathbb{Z}\}$ (LaborieR08). It is fixed if $x=\bot$ (the interval is not executed) or $x=[s,e)$ (the interval starts at $s$ and ends at $e$). For convenience, the description of the filtering algorithm manipulates conditional time-interval domains through the following attributes for each task $i\in T$:
@@
  • $s_{i}\in[\underline{s}_{i},\overline{s}_{i}]$ and $e_{i}\in[\underline{e}_{i},\overline{e}_{i}]$ are the (non negative) start/end variables.
@@
A bound consistent filtering on the relation $s_{i}+d_{i}=e_{i}$ is maintained atomically on the corresponding attributes, and the domain of a time-interval variable becomes $\bot$ whenever one attribute range ($s,d$ or $e$) becomes empty which may trigger an inconsistency if $p_{i}=\mathit{true}$. The whole domain of a conditional time-interval $i\in T$ is thus encoded by a tuple of variables $x_{i}=\langle s_{i},d_{i},e_{i},p_{i}\rangle$.
2.2. Cumulative Functions
For modeling cumulative problems, (LaborieRSV09) introduced the idea of cumulative function expressions. They allow the expression of step-wise integer functions of conditional time-interval variables. The contribution of a single time-interval variable $i\in T$ is represented by an elementary cumulative function receiving an interval $x_{i}$ and a non-negative consumption value $c_{i}$ or a non-negative height range $\underline{c}_{i},\overline{c}_{i}]$. The elementary functions are pulse, stepAtStart and stepAtEnd (LaborieRSV09) as represented in Figure 1.
Figure 1. Elementary cumulative functions.
@@ 2.2. Cumulative Functions
Three tasks $A$, $B$ and $C$ contribute to a cumulative function $f=\mathit{stepAtStart}(A,2)\allowbreak-\allowbreak(pulse(B,1)\allowbreak+\allowbreak\mathit{stepAtEnd}(C,1))$. The resulting AST and cumulative function are represented in Figure 2. On the right, the solid parts represent the time windows of the tasks while the hatched parts represent their contribution to the profile. A dotted line represents the final cumulative function $f$.
Figure 2. Example 1: AST (left) and cumul. function (right).
A cumulative function can be flattened using Algorithm 1 to extract a set of cumulative tasks (tasks with their positive or negative heights represented by a tuple $(x_{i},c_{i})$). This algorithm traverses the AST and collects the activities at the leaf nodes, reversing the height signs for the activities flattened on the right branch of a minus node. Note that whenever encountering a $\mathit{stepAtStart}(x,c)$ leaf node, the cumulative task collected is defined on a fresh task interval $x^{\prime}$ related to the original one with: $x^{\prime}.s=x.s$ (same start as $x$), $x^{\prime}.p=x.p$ (same execution status), and $x^{\prime}.e=Horizon$ (ending at the time horizon) to reflect the definition of a step function. Similarly, in the case of $\mathit{stepAtEnd}(x,c)$, a new task interval $x^{\prime}$ with $x^{\prime}.s=x.e$ (same start as end of $x$) is collected.
Input: $n$: A node of the AST
Output: Set of cumulative tasks
1 if $n=\mathit{pulse}(x,c)$ then $\{(x,c)\}$
2 else if $n=\mathit{stepAtStart}(x,c)$ then $\{(x^{\prime}\leftarrow\langle x.s,Horizon-x.s,Horizon,x.p\rangle,c)\}$
3 else if $n=\mathit{stepAtEnd}(x,c)$ then $\{(x^{\prime}\leftarrow\langle x.e,Horizon-x.e,Horizon,x.p\rangle,c)\}$
4 else if $n=\mathit{plus}(n_{l},n_{r})$ then $\mathit{flatten}(n_{l})\cup\mathit{flatten}(n_{r})$
5 else if $n=\mathit{minus}(n_{l},n_{r})$ then $\mathit{flatten}(n_{l})\cup\{(x,-c)\mid(x,c)\in\mathit{flatten}(n_{r})\}$
6 else errorerror // Unknown pattern
Algorithm 1 flatten($n$)
@@
Continuing example 1, flatten($f$) collects the cumulative tasks $\{(A^{\prime},2),(B,-1),(C^{\prime},-1)\}$ illustrated on the right of Figure 2.
2.3. Constraints on Cumulative Functions
The height of a cumulative function can be restricted with a lower and upper bound at every time overlapping at least one executing time-interval: $\underline{C}\leq f\leq\overline{C}$. This constraint is called $alwaysIn(f,\underline{C},\overline{C})$ in (LaborieRSV09). It can be enforced by posting a GeneralizedCumulative constraint requiring a resource consumption between $\underline{C}$ and $\overline{C}$ for the set of cumulative tasks $T$ collected with the Flatten function. The whole domain of a conditional (cumulative) task $i\in T$ is encoded by a tuple of variables $\langle s_{i},d_{i},e_{i},c_{i},p_{i}\rangle$ where the integer variable $c_{i}$ represents its consumption (possibly negative or hybrid) on the resource. The set of tasks $T$ is partitioned into three subsets as given in Definition 3.
@@ 2.3. Constraints on Cumulative Functions
The set $O=\{i\in T\mid p_{i}=\{\mathit{true},\mathit{false}\}\}$ denotes the set of optional tasks, those for which it is not decided yet if they will execute or not on the resource. The set $R=\{i\in T\mid p_{i}=\mathit{true}\}$ denotes the set of required tasks, those which will surely be executed on the resource. The set $E=\{i\in T\mid p_{i}=\mathit{false}\}$ denotes the set of excluded tasks, those which will surely not be executed on the resource. The sets $O$, $R$ and $E$ are disjoint and form a partition of the set $T$, i.e., $T=O\cup R\cup E$.
@@
The $\texttt{GeneralizedCumulative}(T,\underline{C},\overline{C})$ constraint enforces the cumulated heights of the tasks of $T$ to be within the fixed capacity range $\underline{C},\overline{C}]$ for every time at which at least one task of $T$ is executed:
(1) $\sum\limits_{i\in R\mid s_{i}\leq\tau<e_{i}}c_{i}\in[\underline{C},\overline{C}]\qquad\forall\tau\in\bigcup_{i\in R}[s_{i},e_{i})$
@@
Enforcing global bound-consistency for the cumulative constraint is NP-Hard (GareyJ79). However, enforcing it on the decomposition into sum constraints at every point in the horizon, as in (1), is not NP-Hard. Unfortunately, this approach is inefficient for large horizons, as it requires a sum constraint to be enforced at every time point. Therefore, we aim to achieve the same filtering in polynomial time relative to the number of tasks, rather than the size of the horizon. This filtering is generally referred to as time-tabling filtering. Intuitively, it works by aggregating all the fixed parts (when $\underline{s}_{i}<\overline{e}_{i}$) to create a consumption profile. It then postpones the earliest execution of a task whenever it detects that starting it at this time would overload the resource capacity.
3. Timetabling Algorithm
This section first introduces the Profile data structure then explains the time-tabling filtering algorithm for filtering the task interval attributes.
3.1. The Profile Data Structure
The Profile data structure is an aggregation of adjoined rectangles of different lengths and heights introduced in (GingrasQ16; GayHS15) to record the resource utilization of tasks. The end of a rectangle is the beginning of the next one. The starting (resp. ending) of a rectangle is represented by a tuple called time point with a component representing the beginning (resp. ending) time of the rectangle and other components that contain information relative to the resource consumption for the duration of the rectangle. These time points are sorted in increasing order of time and are kept in a linked list structure called the Profile. An original idea of (GingrasQ16) is to have pointers $tp(\underline{s}_{i})$, $tp(\underline{e}_{i})$ and $tp(\overline{e}_{i})$ linking to the time point associated with $\underline{s}_{i}$, $\underline{e}_{i}$, and $\overline{e}_{i}$ respectively.
3.2. Timetable Filtering
To check and adjust the attributes of tasks, the Profile data structure from (GingrasQ16) is extended to compute both a minimum and maximum profile range at each time point. In our adaptation of the Profile data structure (hereafter referred to as timeline and noted $P$), a time point $tp\in P$ consists of a tuple $\langle\mathit{time},\underline{P},\overline{P},\mathit{\#fp}\rangle$, where $\mathit{time}$ corresponds to the start time of the time point, $\underline{P}$ (resp. $\overline{P}$) to the minimum (resp. maximum) resource profile of tasks over time and $\mathit{\#fp}$ to the number of fixed parts of tasks that overlap over $\mathit{time}$.
@@ 3.2. Timetable Filtering
(2) $\underline{P}_{t}=\sum\limits_{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}\min(\underline{c}_{i},0)+\sum\limits_{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}\max(\underline{c}_{i},0)$
(3) $\overline{P}_{t}=\sum\limits_{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}\max(\overline{c}_{i},0)+\sum\limits_{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}\min(\overline{c}_{i},0)$
@@
3.3. Timeline Initialization and Consistency Check
The function initializeTimeline (Algorithm 2) initializes the timeline, links the corresponding time points to tasks, and computes the minimum and maximum profile of the resource utilization. It receives as input the set of tasks $T$ and the bounds of the resource capacity $\underline{C},\overline{C}]$. First, events are generated by iterating over the tasks that are not absent (Line 2). For each non-absent task $i$, two events are generated which correspond to the minimum start time ($\underline{s}_{i}$), and the maximum end time ($\overline{e}_{i}$) of the task. For tasks that are present and with a fixed part, two additional events are generated for the maximum start time ($\overline{s}_{i}$) and the minimum end time ($\underline{e}_{i}$) of the task.
Three accumulators are used to keep track of the minimum ($\underline{P}$) and maximum ($\overline{P}$) profile as well as the number of fixed parts of tasks overlapping the current time point ($\mathit{\#fp}$). Depending on the nature of the event, the positive or negative contribution of its associated task is either added to (in case of $\underline{s}$ or $\overline{s}$) or subtracted (in case of $\overline{e}$ or $\underline{e}$) from the accumulators and the current time point $tp$ is updated.
As the timeline is initialized, the consistency of the minimum and maximum profiles is checked for each time point at Line 2. Once a time point is completed (when creating the next time point), we ensure that if at least one task is guaranteed to execute during the time point ($\mathit{\#fp}>0$), the minimum profile $\underline{P}$ is not above the maximum capacity of the resource $\overline{C}$ and the maximum profile $\overline{P}$ is not below the minimum capacity of the resource $\underline{C}$.
$\mathit{events}\leftarrow\{\}$
// Initialize the set of events
@@ 3.3. Timeline Initialization and Consistency Check
 $\mathit{events}\leftarrow\mathit{events}\cup\{\underline{s}_{i},\overline{e}_{i}\}$
 // Add events for non-absent tasks
@@
    $\mathit{events}\leftarrow\mathit{events}\cup\{\overline{s}_{i},\underline{e}_{i}\}$
    // Add events for required tasks with fixed part
3
$\underline{P}\leftarrow 0$; $\overline{P}\leftarrow 0$
// Lower/Upper bound of the profile at current time point
$\mathit{\#fp}\leftarrow 0$
// Number of overlapping fixed parts of current time point
4
$tp\leftarrow\textbf{new}\,tp(\min(\mathit{events}.\mathit{time}))$
// Current time point
$P\leftarrow\{tp\}$
// All time points created so-far
5
6for $e\in\mathit{events}$ sorted by time do
@@
8 if $tp.\mathit{time}<e.\mathit{time}$ then
9    if $\mathit{\#fp}>0\wedge(\underline{P}>\overline{C}\vee\overline{P}<\underline{C})$ then
       return Fail
       // Capacity violation detected
10
    $tp\leftarrow\textbf{new}\,tp(e.\mathit{time})$
    // New time point
11    $P.append(tp)$
12
13 if $e=\underline{s}_{i}$ then // Process earliest start event
    $\underline{P}\leftarrow\underline{P}+\min(\underline{c}_{i},0)$
    // Add pessimistic negative contribution
    $\overline{P}\leftarrow\overline{P}+\max(\overline{c}_{i},0)$
    // Add optimistic positive contribution
    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$; $tp(\underline{s}_{i})\leftarrow tp$
    // Update tp and link task to tp
@@
15 if $e=\overline{e}_{i}$ then // Process latest end event
    $\underline{P}\leftarrow\underline{P}-\min(\underline{c}_{i},0)$
    // Remove pessimistic negative contribution
    $\overline{P}\leftarrow\overline{P}-\max(\overline{c}_{i},0)$
    // Remove optimistic positive contribution
    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$; $tp(\overline{e}_{i})\leftarrow tp$
    // Update tp and link task to tp
16
17 if $p_{i}=\mathit{true}\wedge\overline{s}_{i}<\underline{e}_{i}$ then
18    if $e=\overline{s}_{i}$ then // Process start of fixed parts event
       $\underline{P}\leftarrow\underline{P}+\max(\underline{c}_{i},0)$
       // Add pessimistic positive contribution
       $\overline{P}\leftarrow\overline{P}+\min(\overline{c}_{i},0)$
       // Add optimistic negative contribution
       $\mathit{\#fp}\leftarrow\mathit{\#fp}+1$
       // Increment overlapping fixed part counter
19       $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$
20
21    if $e=\underline{e}_{i}$ then // Process end of fixed parts event
       $\underline{P}\leftarrow\underline{P}-\max(\underline{c}_{i},0)$
       // Remove pessimistic positive contribution
       $\overline{P}\leftarrow\overline{P}-\min(\overline{c}_{i},0)$
       // Remove optimistic negative contribution
       $\mathit{\#fp}\leftarrow\mathit{\#fp}-1$
       // Decrement overlapping fixed part counter
22       $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$
@@
// Return the complete timeline
Algorithm 2 initializeTimeline($T,\underline{C},\overline{C}$)
@@
AA:
: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=[1,2],\allowbreak p=\mathit{true}\rangle$;
$B$:
: $\langle s=[2,4],\allowbreak d=[3,4],\allowbreak e=[5,7],\allowbreak c=2,\allowbreak p=\mathit{true}\rangle$;
$C$:
: $\langle s=[3,8],\allowbreak d=[1,3],\allowbreak e=[4,9],\allowbreak c=[-2,1],\allowbreak p=\{\mathit{true},\mathit{false}\}\rangle$;
and a resource $C$ of capacity bounds $[0,1]$. The time points and their attributes are shown in Table 1.
time event(s) $\underline{P}$ $\overline{P}$ $\mathit{\#fp}$